OpenStudy (anonymous):

(sqrt2-sqrt3)(sqrt2+sqrt3) (sqrt2-sqrt3)(sqrt2+sqrt3) @Mathematics

OpenStudy (anonymous):

\[(\sqrt{2}-\sqrt{3)}(\sqrt{2}+\sqrt{3)}\] something like that

OpenStudy (anonymous):

-1

OpenStudy (asnaseer):

use the rule:\[(a-b)(a+b)=a^2-b^2\]

OpenStudy (asnaseer):

in your case:\[a=\sqrt{2}\]\[b=\sqrt{3}\]

OpenStudy (anonymous):

ive worked it a couple of times but i get different answers i dont know what im doing wrong

OpenStudy (asnaseer):

can you show your working please?

OpenStudy (anonymous):

ok give me a min to type it in

OpenStudy (anonymous):

\[\sqrt{4}\]\[ \sqrt{6}\] is that right do far

OpenStudy (anonymous):

\[\sqrt{9} i mean

OpenStudy (asnaseer):

are you trying to multiply out the expression?\[(\sqrt{2}-\sqrt{3})(\sqrt{2}+\sqrt{3})\]

OpenStudy (anonymous):

yea\[\sqrt{4 }+\sqrt{6}-\sqrt{6}-\sqrt{9}\]

OpenStudy (asnaseer):

that is correct, so the sqrt{6}'s will cancel out leaving you:\[\sqrt{4}-\sqrt{9}=2-3=-1\]

OpenStudy (anonymous):

would the 6's cancel out and leave \[\sqrt{4}-\sqrt{9}\]

OpenStudy (asnaseer):

however, if you spot the "pattern" of the question, and if you are familiar with the relation:\[a^2-b^2=(a+b)(a-b)\]then you can do this quicker